OpenStudy (anonymous):

You paint 12 wall in 14 hour. At that rate, how long will it take you to paint one wall?

OpenStudy (chmvijay):

for 12 wall it takes = 14 hr for one wall = 14*1 /12 =

OpenStudy (texaschic101):

or you can set it up as a proportion : 12/14 = 1/x (12 walls to 14 hrs = 1 wall to x hours) cross multiply (12)(x) = (14)(1) 12x = 14 now just solve for x